OpenStudy (anikate):

what is function notation? Using function notation:f(x), g(x), h(x), etc. If f(x)=2x^2 -3x +2, find: f(0), f(2), f(-4) and f(x+3). Again, f(x)=2x^2 -3x +2. Also, suppose that g(x)= x-2. Find f(x)*9(x)

OpenStudy (anikate):

@Calcmathlete

OpenStudy (anonymous):

Is that 9 near the end supposed to be a g?

OpenStudy (anikate):

yes

OpenStudy (anonymous):

And that * sign is a closed dot right? Not a hollow one?

OpenStudy (anikate):

its a multiplying symbol

OpenStudy (anonymous):

Alright. For the f(0), f(2), f(-4), and f(x + 3), just plug them in for x and simplify. For instance, f(0) would be: \[f(0) = 2(0)^2 - 3(0) + 2 \implies 2\] For the f(x) * g(x) part, just multiply them together. \[f(x) * g(x) = (2x^2 - 3x + 2)(x - 2)\]

OpenStudy (anikate):

oh ok for the first one can you just explain f(x+3)

OpenStudy (anonymous):

Same concept. \[f(x + 3) = 2(x + 3)^2 - 3(x + 3) + 2 \implies?\]

OpenStudy (anikate):

ok

OpenStudy (anikate):

plz join my naxt chat question thanks! :D

OpenStudy (anonymous):

Can you hold on for that one? I have my own question to ask.